OpenStudy (anonymous):

It cost $3 to buy a raffle ticket. If there are 1000 tickets sold, and there is one first prize of $500, one second prize of $250 and one third prize of $100, what is the fair price of a ticket?

OpenStudy (anonymous):

so the total amount of award money = 500 + 250 + 100 = $850 i suppose the "fair price" of a ticket would be the price at which there is no leftover money (meaning it's ALL going to the prizes). so this would be $850 / 1000 tickets = $0.85 per ticket.